### v1

# Improving Medical Multi-modal Contrastive Learning with Expert Annotations

Abstract
We introduce eCLIP, an enhanced version of the CLIP model that integrates expert annotations in the form of radiologist eye-gaze heatmaps. It tackles key challenges in contrastive multi-modal medical imaging analysis, notably data scarcity and the “modality gap” – a significant disparity between image and text embeddings that diminishes the quality of representations and hampers cross-modal interoperability. eCLIP integrates a heatmap processor and leverages mixup augmentation to efficiently utilize the scarce expert annotations, thus boosting the model’s learning effectiveness. eCLIP is designed to be generally applicable to any variant of CLIP without requiring any modifications of the core architecture. Through detailed evaluations across several tasks, including zero-shot inference, linear probing, cross-modal retrieval, and Retrieval Augmented Generation (RAG) of radiology reports using a frozen Large Language Model, eCLIP showcases consistent improvements in embedding quality. The outcomes reveal enhanced alignment and uniformity, affirming eCLIP’s capability to harness high-quality annotations for enriched multi-modal analysis in the medical imaging domain.
1 Introduction
Pretraining foundation models on multi-modal data – particularly leveraging the relationships between text and images – has proven to be a robust strategy for generating versatile embeddings (Radford et al., 2021; Jia et al., 2021). These embeddings enhance the efficacy in several downstream tasks, from image generation to advanced vision-language integration (Liu et al., 2023; Ramesh et al., 2021; Singh et al., 2022). Central to this approach is the employment of a contrastive learning (CL) loss objective (Oord et al., 2018; Chen et al., 2020; Zbontar et al., 2021), where models are trained to align positive pairs (e.g., an image and its corresponding caption) while diversifying negative ones. A significant hurdle in this approach is the necessity of vast datasets, often comprising several millions of data points, for competitive results. Models such as CLIP (Radford et al., 2021) have been trained on internet-scale datasets, estimated to encompass hundreds of millions of image-text pairs (Cherti et al., 2023; Xu et al., 2023a). Acquiring datasets of this magnitude poses substantial challenges in specialized fields that require expert knowledge for data collection, processing and annotation. The medical imaging domain exemplifies these difficulties, where acquiring even a single data point, such as a chest X-ray, involves complex processes requiring expertise and significant resources. Moreover, the procurement of such data for machine learning research is further complicated by ethical considerations, patient privacy concerns and the need for extensive de-identification procedures.
This has led to the prevalent use of foundation models, initialized with weights from models trained on extensive internet-scale datasets, for tasks in the medical domain (Zhang et al., 2022; Wang et al., 2022; Huang et al., 2021; Krishnan et al., 2022). However, the areas of interest within medical images are often nuanced and require expert knowledge to interpret, rendering them indistinguishable to a general-purpose model. In 0(a), we investigate the embeddings generated by a CLIP model – initially pretrained on internet data – using samples from the Open-I dataset(Demner-Fushman et al., 2016), which includes X-rays and corresponding radiology reports. We categorize the samples into subgroups based on the primary abnormality identified in each report, such as ‘normal’, cardiomegaly, atelectasis and opacity. A histogram of the cosine similarities between embeddings from different groups indicates a high degree of similarity, with values approaching 1. This could lead to potential challenges in downstream zero-shot inference tasks, which rely on the spatial segregation of embeddings from different groups (Radford et al., 2021). Typically, continual pretraining on medically relevant data is employed to enhance the model’s ability to differentiate between various abnormalities.
((a)) Histogram of Cosine Similarities Among Subgroups. The model exhibits high cosine similarity among embeddings within the text (red) and image (blue) modalities, regardless of the differences among subgroups. This underscores the model’s challenge in capturing the subtleties inherent in medical data.
((b)) Modality Gap. Despite the CLIP contrastive loss aiming at closely aligning image and text embeddings within a shared space, the modalities remain segregated into distinct regions.
Figure 1: Analysis of CLIP Embeddings in Medical Imaging The figure presents embeddings generated by a CLIP model, pretrained on an internet-scale dataset, applied to the Open-I dataset pairing X-rays with corresponding radiology reports.
Recent studies (Liang et al., 2022; Goel et al., 2022; Oh et al., 2023; Zhang et al., 2024; Tschannen et al., 2023) have identified a “modality gap” in multi-modal contrastive representation learning, where the embeddings from different modalities (e.g., images and text) fall in distinct regions in the shared embedding space. This separation, which arises from factors such as initial model weights and the objectives of contrastive learning (Liang et al., 2022; Zhang et al., 2024), leads to the “cone effect” where embeddings of each modality are restricted to a narrow region of the embedding hypersphere. In 0(b), we illustrate this within the medical domain with a 2D UMAP(McInnes et al., 2018) projection of image and text embeddings. This example highlights how embeddings from the same modality but different semantic groups, such as X-ray images of varying abnormalities, cluster closely together. This makes it difficult for a model to distinguish between semantically different images, undermining its performance in medical image analysis.
Figure 2: eCLIP Pretraining with Expert Annotations. eCLIP adds a Heatmap Processor (right), featuring a multi-headed attention layer, to the standard Image and Text encoders in CLIP. This processor, along with vision and text encoders, maps inputs into a shared hypersphere. Here, the original image ($I_{i}$), its text ($T_{i}$) and the heatmap-processed image ($I^{E}_{i}$) are positioned within a tripartite area (shown here after 2D UMAP projection, please refer to the Supplement for a scaled version). We employ mixup between $I_{i}$ and $I^{E}_{i}$ to generate the embedding $v^{\lambda}_{i}$, which gives us additional positive pairs to enhance the CLIP InfoNCE loss optimization. An auxillary loss, $\mathcal{L}_{\text{priming}}$, is used during the initial training steps to “prime” the heatmap processor to imitate an identity function when the heatmap is composed of all ones.
We investigate the potential of integrating expert annotations, specifically radiologist eye-gaze heatmaps, to alleviate these issues. Processing the eye-gaze data from radiologists(Karargyris et al., 2021) provides us with heatmaps indicative of the radiologist’s attention across different regions of the X-ray images. This heatmap reflects areas of clinical interest aligned with details present in radiology reports. We posit that this could help capture nuanced visual cues in the X-rays and therefore pairing it with reports can enrich the CLIP training data with high-quality positive pairs. Due to the scarcity of such expert annotated data, we employ the mixup strategy, a data augmentation technique which has been effective in both supervised (Zhang et al., 2017; Verma et al., 2019; Han et al., 2022) and contrastive learning(Verma et al., 2021; Oh et al., 2023), to create additional synthetic samples.
We present eCLIP (expert-annotated CLIP), an adaptation of the CLIP model that incorporates expert eye-gaze heatmaps, without modifying the CLIP model’s core architecture. The operational workflow of eCLIP is depicted in Fig. 2. Our contributions are as follows:
  • Utilization of Expert Annotations. We harness radiologist eye-gaze heatmaps to create additional embeddings, effectively introducing valuable positive and negative pairs for enhancing the contrastive learning process.
  • eCLIP Architecture. Our implementation features a heatmap processing mechanism utilizing multi-headed attention (MHA), optimized for handling both heatmaps and original images. This is complemented by a mixup strategy to addresses the challenge of data scarcity, and curriculum learning to ensure a gradual introduction of expert annotations.
  • Comprehensive Evaluation. We assess eCLIP’s zero-shot classification accuracy, sample efficiency and cross-modal retrieval performance and embedding quality across multiple chest X-ray datasets. We also evaluate the cross-modal embeddings to generate radiology reports using a frozen Large Language Model (LLM) without explicitly fine-tuning on medical data.
2 Related Work
Modality Gap: Liang et al. (2022) pinpoint the origins of the modality gap to the nuances of model initialization and the objectives of contrastive learning, underscoring its impact on downstream tasks and fairness. Oh et al. (2023) highlight poor uniformity and alignment in CLIP’s embeddings and propose a finetuning method for robust representations. Zhang et al. (2024) explore the geometry of this embedding space, and provide both theoretical and empirical insights on the nature of this geometry. Subsequent research has produced methods to mitigate the modality gap through diverse and creative approaches (Goel et al., 2022; Tschannen et al., 2023; Zhang et al., 2023b; Gu et al., 2023; Nukrai et al., 2022).
Improving Contrastive Learning: Several methods have improved upon the CLIP objective by introducing auxilliary losses, e.g., SLIP (Mu et al., 2022) uses SimCLR (Chen et al., 2020) loss, M3MAE (Geng et al., ; Weers et al., 2023) augment the Masked Autoencoder (He et al., 2022) reconstruction loss, FLIP (Li et al., 2023) randomly masks out input images to improve scaling, DACL (Verma et al., 2021) proposes a domain agnostic mixup strategy, SILC (Naeem et al., 2023) uses self-distillation, Mo et al. (2024) utilized specialist captions to generate pseudo labels for unpaired images, Zhang et al. (2023a) propose Multi-task Paired Masking with Alignment to improve cross-modal interaction. Similarly there have been works that have identified the need to make the CLIP model focus on sub-regions in order to enhance its utility and downstream performance, GLoRIA (Huang et al., 2021) considers the loss from local regions from within the image and reports, Alpha-CLIP (Sun et al., 2023) uses the alpha channel to guide the CLIP model to focus on different regions of the image and generate the masks for all the images in the corpus using an image segmentation pipeline and TIER (Palepu & Beam, 2023) uses a regularization term to improve the local focus of the model.
Multi-modal Contrastive Learning in Medical Imaging: Zhang et al. (2022) demonstrated enhanced downstream performance by jointly using chest X-ray and report pairing for training a contrastive learning model. This was further improved by Huang et al. (2021), by exploiting local and global features from both modalities; Wang et al. (2022) and You et al. (2023) achieved impressive results by using a Swin Tiny model as the image encoder and by adding modifications to the contrastive loss. Several other works have developed similar contrastive learning foundation models while utilizing the biomedical image texts to achieve impressive results (Huang et al., 2023; Xu et al., 2023b; Moor et al., 2023; Wu et al., 2023; Sowrirajan et al., 2010; Tu et al., 2024). Karargyris et al. (2021) and Bigolin Lanfredi et al. (2022) augment a subset of the MIMIC-CXR (Johnson et al., 2019b) samples with high quality eye-tracking and verbal transcripts from several radiologists. van Sonsbeek et al. (2023) and Wang et al. (2024) utilize the heatmaps from eye-gaze to improve image classification.
3 Method
Notations. For a given chest X-ray image $I_{i}$ and its radiology report $T_{i}$, indexed by $i$ in our dataset, we denote their L2-normalized embeddings as $v_{i}$ and $t_{i}$ respectively, residing in a d-dimensional space ($v_{i},t_{i}\in\mathcal{R}^{d}$). Image embeddings are obtained through an encoder $v_{i}=f(I_{i})$ and text embeddings via $t_{i}=g(T_{i})$. Applying an expert heatmap $E_{i}$ to an image results in the corresponding image embedding $v_{i}^{E}$. We denote the loss value for the $i$-th sample as $\mathcal{L}_{i}$. For the case of contrastive loss, this is computed in terms of some similarity measure between the embeddings, $sim(v_{i},t_{i})$, typically cosine similarity defined as $v_{i}\cdot t_{i}$.
3.1 Background
Central to CLIP’s effectiveness is the InfoNCE loss (Oord et al., 2018), a mechanism engineered to optimize the similarity measures between corresponding (positive) pairs and to minimize those among non-corresponding (negative) pairs. The formulation of the CLIP loss objective is as follows:
$\mathcal{L}_{\text{text}}$ $=\mathop{\mathbb{E}}_{(t_{i},v_{i})\sim pos}\left[-\log\frac{\exp(sim(t_{i},v_{i})/\tau)}{\exp(sim(t_{i},v_{i})/\tau)+\sum_{j\neq i}\exp(sim(t_{i},v_{j})/\tau)}\right]$ (1)
Total loss is then defined as, $\mathcal{L}_{\text{total}}=\frac{1}{2}\left(\mathcal{L}_{\text{text}}+\mathcal{L}_{\text{image}}\right)$, where $\mathcal{L}_{\text{image}}$ denotes the corresponding loss for the image to text mapping. Here, $\tau$ represents the temperature parameter that controls the scale of the similarity scores, typically framed as a learnable parameter during training. The loss expectation is taken over all the positive pairings in the dataset.
Theoretical results on CL indicate the concepts of alignment and uniformity as critical for the quality of embeddings (Wang & Isola, 2020; Wang & Liu, 2021). Alignment focuses on reducing the distance between positive pairs while uniformity seeks to evenly distribute the embeddings across the unit hypersphere, preventing extreme clustering that could impair the model’s generalizability and discriminative capabilities. The alignment and uniformity can be defined formally as follows (Oh et al., 2023):
Alignment $=-\mathbb{E}_{(v_{i},t_{i})\sim pos}\left[\|v_{i}-t_{i}\|_{2}^{2}-\min_{j\neq i}\|v_{i}-t_{j}\|_{2}^{2}\right]$ (2)
Uniformity $=-\log\mathbb{E}_{(v_{i},t_{j})\sim\mathcal{D}}\left[\exp(-2\|v_{i}-t_{j}\|_{2}^{2})\right]$ (3)
Figure 3: Comparing eCLIP with $m^{2}$-mixup(Oh et al., 2023). (left) Standard CLIP showing image-text positive pairs $(v_{i},t_{i})$ (solid line), while the other image embeddings serve as negative pairs (dashed line). (center) the $m^{2}$-mixup creates negative pairs $(v^{\lambda}_{j},t_{i})$ via interpolation between embeddings along the geodesic. (right) eCLIP adds expert image embedding, $v^{E}_{i}$, in addition to $v_{i}$ for text $t_{i}$, forming additional positive and negative pairs
High intra-modal similarity, e.g. between two images as seen in Fig. 1, can inadvertently enhance similarity among negative pairs, inflating the denominator of the loss function in Equation (1), and, consequently, hurting the model’s ability to differentiate between positive and negative pairs during training. A conventional method to counter this involves incorporating hard negative pairs, a strategy (Oh et al., 2023) employ by mixing embeddings from different modalities. While effective, this cross-modal mixup may obscure the semantic clarity of embeddings. As an alternative we propose increasing the dataset with additional positive pairs that exhibit minimal semantic overlap by integrating expert annotations. Fig. 3 compares the positive and negative pair creation in eCLIP with traditional CLIP and $m^{2}$-mixup.
3.2 Introducing Expert Annotations to CLIP
Algorithm 1 eCLIP Algorithm
0:  Image Encoder $f(.)$
0:  Text Encoder $g(.)$
1:  $\mathcal{L}_{\text{priming}}\leftarrow 0$; $n_{\text{p}}\leftarrow 0$
2:  for minibatch $\{x_{i}\}_{i=1}^{N}$ do
3:     Unpack $x_{i}$ to $(T_{i},I_{i})$ and optionally $E_{i}$
4:     $t_{i}\leftarrow g(T_{i})$
5:     $v_{i}\leftarrow f(I_{i})$
6:     $p_{\text{uni}}\sim\text{Uniform}(0,1)$
7:     if $p_{\text{uni}}<p_{\text{curr}}$ and $E_{i}$ is provided then
8:        $\lambda\sim\text{Beta}(\alpha,\alpha)$
9:        $I_{i}^{E}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
10:        $I_{i}^{\lambda}\leftarrow I_{i}\lambda+I_{i}^{E}(1-\lambda)$
11:        $v_{i}^{\lambda}\leftarrow f(I_{i}^{\lambda})$
12:     end if
13:     if $E_{i}$ is entirely ones then
14:        $I_{i}^{\text{R}}\leftarrow\text{HeatmapProcessor}(I_{i},E_{i})$
15:        $\mathcal{L}_{\text{priming}}\leftarrow\mathcal{L}_{\text{priming}}+(I_{i}-I_{i}^{\text{R}})^{2}$
16:        $n_{\text{p}}\leftarrow n_{\text{p}}+1$
17:     end if
18:  end for
Our objective with eCLIP is to enhance the CLIP framework by integrating expert annotations – radiologist eye-gaze heatmaps – to diversify the pool of positive samples. The eCLIP model is designed to be compatible across all CLIP variants without modifying its core architecture. The heatmap processor (Fig. 2, right) first converts the images and heatmaps into a sequence of patches and applies multi-headed attention (MHA) over the sequences. The patchified heatmap overlaid images serve as queries, while the original image’s patches act as keys and values. The processed output is then reconstructed back to its original image format, enabling the standard CLIP image encoder to obtain expert image embeddings. These new embeddings and their text embedding pair introduce additional positive samples for the contrastive loss objective (Fig. 3, right).
However, the size of the expert annotated data is orders of magnitude smaller than the data available for CLIP training. To effectively leverage the scarce expert-annotated data, we implement mixup augmentation (Zhang et al., 2017). As illustrated in Fig. 2 (left), this involves blending an original image $I_{i}$ with its expert version $I_{i}^{E}$ to create $I_{i}^{\lambda}=\lambda I_{i}+(1-\lambda)I_{i}^{E}$, where $\lambda\sim\text{Beta}(\alpha,\alpha)$. (We set $\alpha=0.3$ in all our experiments.) The eCLIP image encoder then processes $I_{i}^{\lambda}$ to produce the image embedding $v_{i}^{\lambda}=f(I_{i}^{\lambda})$. These expert embeddings form new positive pairs $(v_{i}^{\lambda},t_{i})$ as well as corresponding negative pairs, which are added with existing pairs $(v_{i},t_{i})$ during the computation of the CLIP InfoNCE Loss, $\mathcal{L}_{i}$.
To seamlessly integrate expert annotations without disrupting the foundational training of the eCLIP model, we employ a phased curriculum learning strategy (Bengio et al., 2009). This approach comprises a cold start phase where the model is initially trained without the expert annotations to establish a robust baseline. This phase accounts for about 10% of the total training iterations. It then transitions into a warmup phase, gradually increasing the inclusion of expert examples from 0.05 to 0.5 probability over the next 30% of iterations. Finally, a cooldown phase reduces expert example probability to 0.1 for the subsequent 40% of iterations, fine-tuning the model’s performance by balancing foundational and expert-driven insights.
Additionally, we regularize the heatmap processor to behave as an identity function in scenarios where the heatmap is entirely composed of ones. We achieve this through a priming phase that coincides with the curriculum learning’s cold start phase. We setup an auxillary mean-squared error loss to force the heatmap processor to reconstruct the original image $I_{i}$ when the heatmap $E_{i}=1$. This priming ensures the heatmap processor’s adaptability, allowing it to process expert annotations effectively when available, while falling back to the model’s original performance in their absence. The total loss during this phase is, $\mathcal{L}_{\text{total}}=w_{\text{p}}\cdot\mathcal{L}_{\text{priming}}+(1-w_{\text{p}})\cdot\mathcal{L}_{\text{clip}}$, where $w_{\text{p}}$ is a hyperparameter which we set to 0.1. The pseudocode for eCLIP is shown in Algorithm 1.
4 Experiments
Our experiments are designed to evaluate the influence of expert heatmap annotations on the quality of the learned representations. Unless stated otherwise, we assume that a large set of image-text pairs, of which a small fraction is annotated with eye-gaze heatmaps, is used for training, but for test samples no annotations are used. We utilize both quantitative measures and qualitative assessments to study the contributions of these annotations towards enhancing model performance.
4.1 Setup
4.1.1 Baselines
[…]
4.1.2 Datasets
For pretraining phase we utilize the MIMIC-CXR dataset (Johnson et al., 2019b), which pairs roughly 200K chest X-rays with free-text radiology reports. The images were processed into JPEG format as described in (Johnson et al., 2019a), and the accompanying reports were stripped of unnecessary punctuation and tokenized using the Wordpiece scheme (Devlin et al., 2018). We obtain the eye-gaze heatmap from the EGD-CXR dataset (Demner-Fushman et al., 2016) and process the eye-tracking data to obtain the normalized eye-gaze heatmap which are available for 1080 datapoints.
Our evaluation setup includes multiple publicly available chest X-ray datasets, specifically CheXpert (Irvin et al., 2019), RSNA Pneumonia (Shih et al., 2019), NIH CXR (Wang et al., 2017) and Open-I (Demner-Fushman et al., 2016), each offering a distinct set of imaging and reporting characteristics. Following previous works (Wang et al., 2022; Huang et al., 2021; You et al., 2023), we prepare the test sets from MIMIC and CheXpert, selecting 200 random samples for five specific pathologies from the CheXpert competition, resulting in 1000 samples for each dataset (MIMIC 5x200 and CheXpert 5x200, respectively). For the NIH-CXR dataset, we assembled a subset of 100 samples for each of 14 abnormalities, thereby creating CXR 14x100 test set. The Open-I dataset is utilized for text retrieval and radiology report generation tasks. For linear probe evaluations, we use CXR-8 (Wang et al., 2017), RSNA dataset and construct an OpenI-5 dataset by extracting labels from the ‘Problems’ field within the reports that match the CheXpert competition labels.
4.1.3 Training
[…]
4.2 Zero-shot Image Classification
Table 1: Zero-shot classification performance on 4 X-ray datasets and model configurations, reported as macro-averaged F1 scores from three independent random seeds. The highest score per dataset and model configuration is underlined. The overall best-performing model for each dataset is highlighted in bold.
Model Dataset
Chexpert 5x200 MIMIC 5x200 RSNA CXR 14x100
[…]
  ${}_{+\text{naive}}$ $0.391_{\pm.069}$ $0.334_{\pm.057}$ $0.731_{\pm.023}$ $0.113_{\pm.024}$
  ${}_{+\text{DACL}}$ $0.506_{\pm.029}$ $0.430_{\pm.011}$ $0.736_{\pm.007}$ $0.158_{\pm.004}$
[…]
  +expert (${}_{+\text{expert (ours)}}$ $0.507_{\pm.004}$ $0.430_{\pm.009}$ $0.761_{\pm.017}$ $0.156_{\pm.023}$
  ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.507_{\pm.005}$ $0.475_{\pm.008}$ $0.775_{\pm.001}$ $0.159_{\pm.001}$
 CLIPSwin ${}_{\text{Swin Tiny}}$ $0.517_{\pm.025}$ $0.452_{\pm.002}$ $0.808_{\pm.000}$ $0.169_{\pm.003}$
  ${}_{+\text{naive}}$ $0.532_{\pm.010}$ $0.452_{\pm.022}$ $0.807_{\pm.007}$ $0.167_{\pm.007}$
  ${}_{+\text{DACL}}$ $0.465_{\pm.008}$ $0.389_{\pm.015}$ $0.768_{\pm.018}$ $0.101_{\pm.013}$
[…]
  +expert (${}_{+\text{expert (ours)}}$ $0.549_{\pm.016}$ $0.445_{\pm.021}$ $0.818_{\pm.004}$ $0.172_{\pm.006}$
  ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.558_{\pm.004}$ $0.463_{\pm.007}$ $0.819_{\pm.000}$ $0.192_{\pm.003}$
 CLIPViT ${}_{\text{ViT Small}}$ $0.525_{\pm.024}$ $0.441_{\pm.006}$ $0.807_{\pm.006}$ $0.159_{\pm.007}$
  ${}_{+\text{naive}}$ $0.534_{\pm.016}$ $0.440_{\pm.019}$ $0.805_{\pm.004}$ $0.156_{\pm.017}$
  ${}_{+\text{DACL}}$ $0.475_{\pm.025}$ $0.398_{\pm 015}$ $0.761_{\pm.009}$ $0.133_{\pm.007}$
[…]
  +expert (${}_{+\text{expert (ours)}}$ $0.545_{\pm.016}$ $0.452_{\pm.013}$ $0.803_{\pm.003}$ $0.165_{\pm.007}$
  ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.559_{\pm.001}$ $0.439_{\pm.004}$ $0.817_{\pm.001}$ $0.165_{\pm.004}$
 CLIPViT ${}_{\text{ViT Base}}$ $0.540_{\pm.017}$ $0.465_{\pm.004}$ $0.805_{\pm.001}$ $0.183_{\pm.011}$
  ${}_{+\text{naive}}$ $0.506_{\pm.011}$ $0.426_{\pm.006}$ $0.805_{\pm.004}$ $0.151_{\pm.009}$
  ${}_{+\text{DACL}}$ $0.474_{\pm.007}$ $0.400_{\pm.002}$ $0.759_{\pm.001}$ $0.106_{\pm.003}$
[…]
  +expert (${}_{+\text{expert (ours)}}$ $0.563_{\pm.021}$ $0.477_{\pm.004}$ $0.814_{\pm.003}$ $0.193_{\pm.017}$
  ${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ $0.549_{\pm.003}$ $0.452_{\pm.007}$ $0.810_{\pm.001}$ $0.185_{\pm.002}$
Following CLIP(Radford et al., 2021), our zero-shot classification method categorizes images into predefined classes without direct finetuning, thus relying on the quality of embeddings generated during pretraining. To formulate the embedding of each class label, we first generate descriptive prompts to obtain a list of text embeddings corresponding to the label using the text encoder (Huang et al., 2021; Wang et al., 2022; You et al., 2023). The mean of these embeddings is taken as the representation of the label. For each image, classification is then performed by matching the image embedding with its closest label embedding through cosine similarity. More details of the prompts used are provided in the Supplement.
Table 1 illustrates the zero-shot classification performance on CheXpert 5x200, MIMIC 5x200, RSNA, and CXR 14x100 datasets. The results, based on macro-averaged F1 scores from three random initializations, highlight eCLIP variants’ superior performance over the base models across all datasets. While the $m^{3}$-mixup excels in MIMIC for certain architectures, eCLIP variants show broader generalization. Notably, eCLIP’s advantages are more pronounced in multi-class scenarios (CheXpert, MIMIC and CXR14) compared to binary classification on RSNA.
4.3 Sample Efficiency
Figure 4: Sample Efficiency. (top row) Zero-shot performance on three multi-label classification test sets for CLIP and eCLIP Swin Tiny models, trained with varying amounts of training batches. (bottom row) Linear probe scores with varying amounts of training data.
Sample efficiency measures how well a model learns from limited amount of training data. eCLIP improves this efficiency by using expert annotated images to form new positive and negative pairs, aiming to improve the quality of the learned embeddings. To test this, we first looked at zero-shot classification performance, adjusting the number of training batches available for pretraining. Results, shown in Fig. 4 (top row), reveal that eCLIP is more sample efficient across MIMIC 5x200, CheXpert 5x200 and CXR 14x100 datasets compared to the base model. Additionally, by applying supervised fine-tuning (SFT) with a linear probe on class-imbalanced datasets – CXR-8, RSNA and OpenI-5 (Section 4.1.2) – eCLIP demonstrates stronger performance in multi-label classification tasks, CXR-8 and OpenI-5 and remains competitive in binary classification for RSNA. This is shown Fig. 4 (bottom row) where we plot the ROC AUC scores against different training sample sizes for linear probing. These findings highlight eCLIP’s ability to effectively learn from fewer samples.
4.4 Text retrieval and retrieval augmented generation (RAG)
Table 2: Performance on Image to Report retrieval with Open-I dataset. We report the Recall@{1, 5, 10}
[…]
Table 4: Ablation Study with Swin Tiny. Zero-shot performance on CheXpert (CXP) and CXR14 (C14) datasets for the base CLIP and models with expert annotation integration (+E) is presented. Methods include mask multiplication ($\odot$), CNN, and Multi-headed Attention (MHA) encoders. Key augmentations: Mixup (+M), Curriculum Learning (+C), and Encoder Priming (+P) demonstrate performance gains. A control with a random mask (rand) confirms the significance of expert annotations. We report macro-averaged F1 scores from three random initialization
Model R@1 R@5 R@10
CLIPSwin ${}_{\text{Swin Tiny}}$ 3.1 7.6 11.3
[…]
+expert (${}_{+\text{expert (ours)}}$ 3.7 9.4 13.4
${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ 3.1 8.2 11.7
CLIPViT ${}_{\text{ViT Base}}$ 3.7 9.2 13.2
[…]
+expert (${}_{+\text{expert (ours)}}$ 4.4 10.3 13.5
${}_{+\text{expert}^{\mathcal{P}}\text{(ours)}}$ 2.9 7.5 10.3
[…]
 CLIP 0.172 0.713 0.791 0.492
 eCLIP 0.177 0.712 0.795 0.506
Method CXP C14
Base $0.517_{\pm.024}$ $0.169_{\pm.003}$
$\odot$ Mask (+ E) $0.540_{\pm.019}$ $0.165_{\pm.006}$
CNN Encoder (+ E) $0.534_{\pm.012}$ $0.163_{\pm.008}$
MHA Encoder (+ E) $0.534_{\pm.013}$ $0.153_{\pm.002}$
MHA Encoder
(+ E, M)
$0.532_{\pm.018}$ $0.160_{\pm.010}$
MHA Encoder
(+ E, M, C)
$0.545_{\pm.008}$ $0.173_{\pm.018}$
MHA Encoder
(+ rand, M, C, P)
$0.537_{\pm.020}$ $0.166_{\pm.003}$
MHA Encoder
(+ E, M, C, P)
$0.549_{\pm.016}$ $0.172_{\pm.006}$
[…]
Table 4: Ablation Study with Swin Tiny. Zero-shot performance on CheXpert (CXP) and CXR14 (C14) datasets for the base CLIP and models with expert annotation integration (+E) is presented. Methods include mask multiplication ($\odot$), CNN, and Multi-headed Attention (MHA) encoders. Key augmentations: Mixup (+M), Curriculum Learning (+C), and Encoder Priming (+P) demonstrate performance gains. A control with a random mask (rand) confirms the significance of expert annotations. We report macro-averaged F1 scores from three random initialization
Table 5: Random samples of generated report. For each image in the Open-I dataset, the five closest text snippets based on embedding cosine similarity is used as prompts for Mistral 7B LLM. Utilizing in-context learning, we prompt the LLM with two such snippet-report pairs. The conditions that the generated report identified correctly are highlighted in green while those it missed are shown in red.
Problems Ground Truth Generated
normal The lungs and pleural spaces show no acute abnormality. Heart size and pulmonary vascularity within normal limits. The lungs are clear with no evidence of consolidation, pneumothorax, or pleural effusion. Heart size and mediastinal contour are normal. No acute abnormalities noted in the visualized osseous structures.
Pulmonary Atelectasis. Pleural Effusion. Cardiomegaly. Pulmonary Congestion AP and lateral views were obtained. Bibasilar atelectasis and small left-sided pleural effusion. Stable cardiomegaly. No pneumothorax. Mild pulmonary vascular congestion. The lungs show bibasilar airspace opacities, possibly due to infection or atelectasis. Heart size is mildly enlarged. No pneumothorax or pleural effusion.
Lung. Density. Emphysema There is hyperexpansion. The heart size is normal. There is no pleural effusion or pneumothorax. Two circular densities overlying the right ribs which were not present in the XXXX CT. No focal infiltrates The lungs are hyperexpanded with clear fields and no acute consolidations or pleural effusions. Heart size and pulmonary vascularity are within normal limits. Degenerative changes are present in the thoracic spine. A calcified granuloma is noted in the left perihilar region. No pneumothorax or large pleural effusion is observed.
To compare eCLIP’s cross-modal functionality with that of CLIP we focused on text retrieval task using the Open-I dataset, which consists of pairs of X-rays and radiology reports. We used the FAISS vector database (Douze et al., 2024) to index the text embeddings generated by the text encoder. For a given X-ray image $I_{i}$, we then retrieve the closest text reports from the database based on the cosine similarity in the embedding space, $min_{j}(v_{i}\cdot t_{j})$. Results in Table 4 compare the performance of eCLIP against CLIP in text retrieval measured in Recall@1, 5, and 10. The performance of eCLIP indicates a notable improvement in its embedding quality. Note that our evaluation followed a strict criterion for recall computation, where a retrieval was counted as successful only if the exact correct report was identified. While more nuanced measures based on semantic similarity could be employed (You et al., 2023), we opted this approach to maintain a clear and simple evaluation framework.
Next we extend our analysis from retrieval to report generation using a frozen Large Language Model (LLM), Mistral 7B Instruct v2 (Jiang et al., 2023), aiming to generate radiology reports through Retrieval Augmented Generation (RAG). This setup tests the CLIP model’s capacity to retrieve texts which can be used to prompt an LLM to generate a report without finetuning on medical data. First we randomly selected 389 samples from the Open-I dataset for testing and utilized the FAISS database to index the reports from the remaining samples (i.e., training set). Given a test image we retrieve five closest reports from the training set and use them in the prompt for the LLM to generate a report for the test image. The eCLIP variant showed a small but consistent improvement over the base model in generating reports, as indicated in Table 4. A comparative analysis of generated report versus ground truth shown in Table 5, with discrepancies marked, further validates the effectiveness of eCLIP’s embeddings in supporting complex cross-modal tasks. Additional details, including LLM prompts and generated report samples are available in the Supplement.
4.5 Embedding Quality
Figure 5: Qualitative Analysis of CLIP Pretraining. top row illustrates the cosine similarity distributions for CLIP and eCLIP image embeddings. bottom left and center sections display uniformity, alignment, and modality gap comparisons among the internet pretrained model (ZS), CLIP pretrained on MIMIC, and eCLIP. bottom right details K-means clustering metrics for image embeddings with k=5 for both CLIP and eCLIP models.
For qualitative evaluations, we first examine the histogram of the cosine similarities of the embeddings from different abnormality subgroups obtained from the CLIP image encoder. In Fig. 5 (top row), we can see that the similarities for the CLIP model has considerably dropped below 1 after continual pretraining on MIMIC-CXR compared to Fig. 1. This indicates that the model’s ability to distinguish between different conditions has improved. The introduction of expert annotations in the eCLIP variant further improves this with mean cosine similarities for ‘normal’ versus cardiomegaly, atelectasis and opacity dropping to 0.36, 0.4 and 0.4 respectively.
Our evaluation of uniformity and alignment reveals that eCLIP surpasses both the MIMIC-pretrained and internet-pretrained models in these key metrics, indicating a marked improvement in the quality of embeddings (Fig. 5, bottom left). We also note a modest decrease in the modality gap with eCLIP (Fig. 5, bottom center). Clustering analysis via K-means (with k=5 for 5 abnormalities in data) highlights eCLIP’s superior performance in grouping abnormalities, as seen from improved scored in Normalized Mutual Information (NMI), Silhouette score, and Calinski-Harabasz (CH) index (Fig. 5, bottom right).
4.6 Ablation Study
Our ablation study with the Swin Tiny encoder shows the impact of key components in our eCLIP model: multi-headed attention (MHA) layer for heatmap procesing, curriculum learning for phased introduction of expert annotations, mixup augmentation to compensate for limited number of export annotated data and priming of heatmap processor during initial training phase. Results shown in Table 4 reveal that the MHA-based heatmap processor improves zero-shot classification performance on CheXpert 5x200 and CXR 14x100 datasets compared to basic methods like direct application of heatmaps as mask ($\odot Mask)$ or using a CNN encoder. We note a significant performance drop with randomly generated heatmaps versus expert eye-gaze heatmaps. This highlights that while our methodological improvements contribute to the performance gains, the integration of meaningful, expert-derived signals is essential for achieving optimal results.
5 Conclusion
We introduce eCLIP, an adaptation of CLIP, demonstrating the integration of radiologist eye-tracking heatmap to overcome challenges faced in multi-modal contrastive learning. This study highlights the impact of integrating these high-quality expert annotations on improving the quality of learned embeddings and assess its influence on sample efficiency and cross-modal retrieval tasks. An important future research direction would be extending this approach to include expert annotations from the text modality (e.g., by adapting SimCSE (Gao et al., 2021)) and to leverage the temporal dynamics of eye-tracking data by aligning the sequential frames with the corresponding report snippets.
5.0.1 Limitations
Our study is limited by the small size of expert annotated data and thus does not comprehensively analyze the impact of size or distribution of expert annotations across different abnormalities. eCLIP also incurs extra computational costs during training due to the additional forward pass required for processing expert images in the warmup and cool-down phases. Additionally, the clinical relevance of generated radiology reports has not been validated by medical experts, relying instead on standard metrics known for potential biases and inaccuracies in reflecting clinical accuracy(Yu et al., 2023).